During a particular baseball season, a team averaged 41,602 fans per home game. Suppose attendance during the season follows the normal probability distribution with a standard deviation of 9,440 per game. What is the probability that a randomly selected game during the season had an attendance greater than 50,000 people?

A.

0.1867

B.

0.0417

C.

0.3255

D.

0.3986
Answer #1

P(X < A) = P(Z < (A - mean)/standard deviation)

P(attendace greater than 50,000) = 1 - P(X < 50,000)

= 1 - P(Z < (50,000 - 41602)/9,440)

= 1 - P(Z < 0.89)

= 1 - 0.8133

= 0.1867